# Compactly supported $\mathbb{A}^{1}$-Euler characteristics of symmetric powers of cellular varieties

### v1

@@ 6. Curves of genus $\leq 1$
In this section we show that curves of geometric genus $\leq 1$ are symmetrisable, but curves of geometric genus $>0$ are not $\operatorname{K}_{0}$-étale linear. Therefore, curves of genus $1$ give examples to show the inclusion $\operatorname{K}_{0}(\mathrm{\acute{E}tLin}_{k})\subseteq\operatorname{K}_{0}(\mathrm{Sym}_{k})$ is strict.
6.1. Curves of genus $\leq 1$ are symmetrisable
@@
Let $C$ be a smooth projective curve, and let $n$ be odd. Then $\chi^{\mathrm{mot}}(C^{(n)})=a_{n}(\chi^{\mathrm{mot}}(C))$.
@@
Note that $C^{(n)}$ is smooth projective of odd dimension, so $\chi^{\mathrm{mot}}(C^{(n)})$ is hyperbolic by [Lev20, Corollary 3.2]. Similarly, since $\chi^{\mathrm{mot}}(C)$ is hyperbolic and $n$ is odd, $a_{n}(\chi^{\mathrm{mot}}(C))$ is hyperbolic by Lemma 2.10, so we only need to check that these have the same rank. Since the rank is invariant under base change of fields, we may argue as in the proof of [PP22, Theorem 2.10] to reduce the result to the case where $k=\mathbb{C}$, where the result holds by a result of MacDonald [Mac62a, (4.4)]. ∎
@@
Smooth projective curves of genus 00 are symmetrisable.
@@
Let $C$ be a smooth projective curve of genus 00. By the above proposition, we only need to show $\chi^{\mathrm{mot}}(C^{(n)})=a_{n}(\chi^{\mathrm{mot}}(C))$ when $n$ is even. If $C=\mathbb{P}^{1}$, then $C$ is linear, so the result follows by Theorem 4.10. Assume therefore that $C$ is a conic with $C(k)=\emptyset$, so there exists a quadratic extension $L/k$ with $C(L)\neq\emptyset$. In particular, we see there is a closed embedding $\operatorname{Spec}(L)\hookrightarrow C$. We then see that $\operatorname{Spec}(L)^{(n)}$ is a closed subvariety of $C^{(n)}$. Moreover, since $C^{(n)}$ is a $k$-form of $\mathbb{P}^{1}$, we see that $C^{(n)}$ will be given by a $k$-form of $(\mathbb{P}^{1})^{(n)}=\mathbb{P}^{n}$, so is also a Severi–Brauer variety. Note that $\operatorname{Spec}(L)^{(n)}$ is given by the disjoint union of $\operatorname{Spec}(k)$ and $\frac{n}{2}$ copies of $\operatorname{Spec}(L)$, so in particular, $\operatorname{Spec}(k)$ embeds as a closed subvariety into $C^{(n)}$, so $C^{(n)}(k)\neq\emptyset$. Since $C^{(n)}$ is a Severi–Brauer variety however, we see $C^{(n)}=\mathbb{P}^{n}$ again by Châtelet’s Theorem ([Poo17, Proposition 4.5.10]). The result is now clear. ∎
@@
Let $C$ be a smooth projective curve of genus $g>0$. Suppose that $n>2g-2$. Then $\chi^{\mathrm{mot}}(C^{(n)})=0$.
@@
Following the proof of [Mus11, Theorem 7.33], we may define a morphism $C^{(n)}\to\text{Pic}^{n}(C)$ which is defined on points over the algebraic closure $\bar{k}$ of $k$ as follows. A point of $C^{(n)}$ over $\bar{k}$ is a divisor $D$ on $C$ of degree $n$ and we send $D$ to $\mathcal{O}_{C}(D)$. If $n>2g-2$, this map makes $C^{(n)}$ into a Zariski locally trivial bundle of degree $n-g$ over $\text{Pic}^{n}(C)$ whose fibre $B^{n-g}$ is a Severi–Brauer variety of dimension $n-g$. We now find that
$\chi^{\mathrm{mot}}(C^{(n)})$ $=\chi^{\mathrm{mot}}(B^{n-g})\chi^{\mathrm{mot}}(\text{Pic}^{n}(C)).$
We note that $\chi^{\mathrm{mot}}(\text{Pic}^{n}(C))=0$ by Theorem 5.48 of [PP22]. This gives the desired result. ∎
@@
We first reduce to the case where $C$ is smooth and projective. Let $\overline{C}$ be a compactification of $C$, and let $\tilde{C}$ be a normalisation of $\overline{C}$. Then $\tilde{C}$ and $C$ are birational and dimension $1$, so the rational map $\tilde{C}\dashrightarrow C$ allows us to realise $[\tilde{C}]-[A]=[C]-[A^{\prime}]$, where $A,A^{\prime}$ are dimension 00 varieties. Since $\operatorname{K}_{0}(\mathrm{Sym}_{k})$ is an sub-abelian group of $\operatorname{K}_{0}(\mathrm{Var}_{k})$ and dimension 00 varieties are symmetrisable by [PP23, Corollary 4.30], $[C]$ is an element of $\operatorname{K}_{0}(\mathrm{Sym}_{k})$ if and only if $[\tilde{C}]$ is. Therefore without loss of generality, assume $C$ is smooth and projective.
If $g(C)=0$, we appeal to Corollary 6.2. If $g(C)=1$, Proposition 6.3 guarantees that $\chi^{\mathrm{mot}}(C^{(n)})=a_{n}(\chi^{\mathrm{mot}}(C))$ for all $n>0$. Therefore, for all $n$, we have that $\chi^{\mathrm{mot}}(C^{(n)})=a_{n}(\chi^{\mathrm{mot}}(C))$, so $[C]\in\operatorname{K}_{0}(\mathrm{Sym}_{k})$. ∎
6.2. Curves of genus $>0$ are not $\operatorname{K}_{0}$-étale linear
Theorem 6.5.
Let $X/k$ be a geometrically connected variety which is not geometrically stably rational. Then $X$ is not $\operatorname{K}_{0}$-étale linear.
@@ 6.2. Curves of genus $>0$ are not $\operatorname{K}_{0}$-étale linear
Let $\overline{k}$ be an algebraic closure of $k$. Suppose that $[X]\in\operatorname{K}_{0}(\mathrm{\acute{E}tLin}_{k})$. Then clearly after base changing to $\overline{k}$, we have $X_{\overline{k}}\in\operatorname{K}_{0}(\mathrm{\acute{E}tLin}_{\overline{k}})$. Since $\overline{k}$ is algebraically closed, all varieties of dimension 00 are disjoint unions of $\operatorname{Spec}(\overline{k})$, and therefore $\operatorname{K}_{0}(\mathrm{\acute{E}tLin}_{\overline{k}})$ is simply the subring of $\operatorname{K}_{0}(\mathrm{Var}_{k})$ generated by $\mathbb{A}^{1}$. This question therefore reduces to the case where $k$ is algebraically closed.
Let $I$ denote the ideal of $\operatorname{K}_{0}(\mathrm{Var}_{k})$ generated by $\mathbb{A}^{1}$. Then by a result of Larsen and Lunts [LL03, Proposition 2.7], we have an isomorphism $\operatorname{K}_{0}(\mathrm{Var}_{k})/I\cong\mathbb{Z}[SB]$, where $\mathbb{Z}[SB]$ is the ring whose underlying additive group is the free abelian group generated by stable-birational classes of varieties. By [LL03, Proposition 2.7], a connected smooth projective variety $X/k$ is stably rational if and only if the class $[X]$ in $\operatorname{K}_{0}(\mathrm{Var}_{k})$ is congruent to $1\pmod{I}$. By assumption, $X$ is not stably rational and is connected, so $[X]\neq 1\in\mathbb{Z}[SB]$, which implies that $[X]\neq n\in\mathbb{Z}[SB]$ for some $n\in\mathbb{Z}$, since $X$ is connected. This is precisely saying that there is no element $Y\in I$ such that $Y+n[\operatorname{Spec}(k)]=[X]$ for any $n\in\mathbb{Z}$. Since every element of $\operatorname{K}_{0}(\mathrm{\acute{E}tLin}_{\overline{k}})$ can be written in this manner, $[X]$ is not in $\operatorname{K}_{0}(\mathrm{\acute{E}tLin}_{\overline{k}})$. ∎
Corollary 6.6.
Let $X$ be a geometrically connected curve of geometric genus $>0$. Then $X$ is not $\operatorname{K}_{0}$-étale linear.
Proof.
As above, we may reduce to the case where $k$ is algebraically closed and $X$ is connected, and as in Corollary 6.4, we may reduce to the case where $X$ is smooth and projective. Suppose $X$ is stably rational, so it is unirational. Since $X$ is a curve, this would imply $X$ is rational by a result of Luroth ([Lür75]). By the Riemann–Hurwitz formula $g(X)=0$. For $X$ not genus 00, it is not geometrically stably rational, so we can apply the above theorem. ∎
As mentioned in Remark 4.15, the above shows that the inclusion $\operatorname{K}_{0}(\mathrm{\acute{E}tLin}_{k})\subseteq\operatorname{K}_{0}(\mathrm{Sym}_{k})$ is always strict. In [BV24], it is shown that all curves are symmetrisable. Since we know that $\operatorname{K}_{0}(\mathrm{Sym}_{k})$ is a module over $\operatorname{K}_{0}(\mathrm{\acute{E}tLin}_{k})$, we can combine the above result with the result from [BV24] to obtain large classes of symmetrisable varieties which are not $\operatorname{K}_{0}$-étale linear.
@@
For other invariants in motivic homotopy theory, we may obtain obstructions to nice behaviour for varieties which are not $\mathbb{A}^{1}$-connected. For example, [KLSW23b, Definition 2.30] introduces the notion of a global $\mathbb{A}^{1}$-degree of a morphism $f:X\to Y$, which descends to an element of $\operatorname{GW}(k)$ only when $Y$ is $\mathbb{A}^{1}$-connected. Therefore, one might expect $\mathbb{A}^{1}$-connectedness to be a necessary condition for symmetrisability. However, when $k$ is a number field, there exist smooth projective curves of genus $\geq 1$ that are not $\mathbb{A}^{1}$-connected. It is therefore not true that over a number field $\operatorname{K}_{0}(\mathrm{Sym}_{k})$ is given by the Grothendieck group of $\mathbb{A}^{1}$-connected varieties.

### v2

@@ 6. Curves of genus $\leq 1$
In this section we show that curves of geometric genus $\leq 1$ are symmetrisable, but curves of geometric genus >0>0 are not $\mathrm{K}_{0}$-étale linear in characteristic 0. Therefore, curves of genus $1$ give examples to show the inclusion $\mathrm{K}_{0}(\mathrm{\acute{E}tLin}_{k})\subseteq\mathrm{K}_{0}(\mathrm{Sym}_{k})$ is strict.
Remark 6.1.
The following proposition was initially proven only in characteristic 0 and before the paper [BV25], which shows that all curves are symmetrisable, both in characteristic 0 and in characteristic >2>2. We keep the below argument in this paper for completeness, however, we use an argument from [BV25] to extend our original argument to characteristic $p$.
@@
Let $C$ be a smooth projective curve, and let $n$ be odd. Then $\chi_{c}^{\mathrm{mot}}(C^{(n)})=a_{n}(\chi_{c}^{\mathrm{mot}}(C))$.
@@
Note that $C^{(n)}$ is smooth projective of odd dimension, so $\chi_{c}^{\mathrm{mot}}(C^{(n)})$ is hyperbolic by [Lev20, Corollary 3.2]. Since the rank is invariant under base change of fields, we may assume $k$ is algebraically closed by base changing to the algebraic closure of $k$.
If $\mathrm{char}(k)=0$ we may argue as in the proof of [PPar, Theorem 2.14] to reduce the result to the case where $k=\mathbb{C}$, where the result holds by a result of MacDonald [Mac62a, (4.4)]. In positive characteristic, we may argue as in [BV25, Lemma 9, Proposition 10] to reduce to the case where $k$ has characteristic 0.
@@
Smooth projective curves of genus 0 are symmetrisable.
@@
Let $C$ be a smooth projective curve of genus 0. By the above proposition, we only need to show $\chi_{c}^{\mathrm{mot}}(C^{(n)})=a_{n}(\chi_{c}^{\mathrm{mot}}(C))$ when $n$ is even. If $C=\mathbb{P}^{1}$, then $C$ is linear, so the result follows by Theorem 4.10. Assume therefore that $C$ is a conic with $C(k)=\emptyset$, so there exists a quadratic extension $L/k$ with $C(L)\neq\emptyset$. In particular, we see there is a closed embedding $\operatorname{Spec}(L)\hookrightarrow C$. We then see that $\operatorname{Spec}(L)^{(n)}$ is a closed subvariety of $C^{(n)}$. Moreover, since $C^{(n)}$ is a $k$-form of $\mathbb{P}^{1}$, we see that $C^{(n)}$ will be given by a $k$-form of $(\mathbb{P}^{1})^{(n)}=\mathbb{P}^{n}$, so is also a Severi–Brauer variety. Note that $\operatorname{Spec}(L)^{(n)}$ is given by the disjoint union of $\operatorname{Spec}(k)$ and $\frac{n}{2}$ copies of $\operatorname{Spec}(L)$, so in particular, $\operatorname{Spec}(k)$ embeds as a closed subvariety into $C^{(n)}$, so $C^{(n)}(k)\neq\emptyset$. Since $C^{(n)}$ is a Severi–Brauer variety however, we see $C^{(n)}=\mathbb{P}^{n}$ again by Châtelet’s Theorem ([Poo17, Proposition 4.5.10]). The result is now clear. ∎
@@
Let $C$ be a smooth projective curve of genus $g>0$. Suppose that n>2g2n>2g-2. Then $\chi_{c}^{\mathrm{mot}}(C^{(n)})=0$.
@@
Let $k^{p^{-\infty}}$ denote the perfect closure of $k$, which is the field given by adjoining all $p^{n}$ roots of elements in $k$ for every $n$ (see [Gre65, 2.]). Since we are not in characteristic $2$, the map $\mathrm{GW}(k)\to\operatorname{GW}(k^{perf})$ is an isomorphism, so without loss of generality, by base changing to $k^{perf}$, we may assume our base field is perfect.
As in the proof of [Mus11, Theorem 7.33], define a morphism $C^{(n)}\to\text{Pic}^{n}(C)$ which is defined on points over the algebraic closure $\bar{k}$ of $k$ as follows. A point of $C^{(n)}$ over $\bar{k}$ is a divisor $D$ on $C$ of degree $n$ and we send $D$ to $\mathcal{O}_{C}(D)$. If n>2g2n>2g-2, this map makes $C^{(n)}$ into a Zariski locally trivial bundle of degree ngn-g over $\text{Pic}^{n}(C)$ whose fibre $B^{n-g}$ is a Severi–Brauer variety of dimension ngn-g. We now find that
$\chi_{c}^{\mathrm{mot}}(C^{(n)})$ $=\chi_{c}^{\mathrm{mot}}(B^{n-g})\chi_{c}^{\mathrm{mot}}(\text{Pic}^{n}(C)).$
Since $\chi_{c}^{\mathrm{mot}}(\text{Pic}^{n}(C))=0$ by [PPar, Theorem 5.29], this gives the result. The result [PPar, Theorem 5.29] is stated in characteristic 0, but an identical argument holds over perfect fields of characteristic $\neq 2$. ∎
@@
We first reduce to the case where $C$ is smooth and projective. Let $\overline{C}$ be a compactification of $C$, and let $\tilde{C}$ be a normalisation of $\overline{C}$, so that $\tilde{C}$ is smooth and projective. Then $\tilde{C}$ and $C$ are birational and dimension $1$, so the rational map $\tilde{C}\dashrightarrow C$ allows us to realise $\tilde{C}]-[A]=[C]-[A^{\prime}]$, where $A,A^{\prime}$ are dimension 0 varieties. Since $\mathrm{K}_{0}(\mathrm{Sym}_{k})$ is an sub-abelian group of $\mathrm{K}_{0}(\mathrm{Var}_{k})$ and dimension 0 varieties are symmetrisable by [PP25, Corollary 4.30], $[C]$ is an element of $\mathrm{K}_{0}(\mathrm{Sym}_{k})$ if and only if $\tilde{C}$] is. Therefore, replacing $C$ with $\tilde{C}$ if necessary, assume $C$ is smooth and projective.
If $g(C)=0$, we appeal to Corollary 6.3. If $g(C)=1$, Proposition 6.4 guarantees that $\chi_{c}^{\mathrm{mot}}(C^{(n)})=0$ for all $n>0$. Since $C^{(1)}=C$, the result for $n=1$ gives us
$\chi_{c}^{\mathrm{mot}}(C^{(n)})=0=a_{n}(0)=a_{n}(\chi_{c}^{\mathrm{mot}}(C)),$
so $C]\in\mathrm{K}_{0}(\mathrm{Sym}_{k})$ as required. ∎
Lemma 6.6.
Let $X$ be a geometrically connected curve of geometric genus >0>0, and suppose $\mathrm{char}(k)=0$. Then $X$ is not $\mathrm{K}_{0}$-étale linear.
@@
As above, we may reduce to the case where $k$ is algebraically closed and $X$ is connected, and as in Corollary 6.5, we may reduce to the case where $X$ is smooth and projective. Suppose $X$ is stably rational, so it is unirational. Since $X$ is a curve, this would imply $X$ is rational by a result of Luroth ([Lür75]). By the Riemann–Hurwitz formula $g(X)=0$. For $X$ not genus 0, it is not geometrically stably rational, so we can apply Theorem 3.3 to get the result. ∎
As mentioned in Remark 4.14, this shows that the inclusion $\mathrm{K}_{0}(\mathrm{\acute{E}tLin}_{k})\subseteq\mathrm{K}_{0}(\mathrm{Sym}_{k})$ is always strict in characteristic 0. In [BV25, Proposition 26], it is shown that all curves are symmetrisable. Since $\mathrm{K}_{0}(\mathrm{Sym}_{k})$ is a module over $\mathrm{K}_{0}(\mathrm{\acute{E}tLin}_{k})$, we use this result to obtain large classes of symmetrisable varieties which are not $\mathrm{K}_{0}$-étale linear, for example, the product of a curve and any $\mathrm{K}_{0}$-étale linear variety.
@@
For other invariants in motivic homotopy theory, we may obtain obstructions to nice behaviour for varieties which are not $\mathbb{A}^{1}$-connected. For example, [KLSW23b, Definition 2.30] introduces the notion of a global $\mathbb{A}^{1}$-degree of a morphism $f:X\to Y$, which descends to an element of $\operatorname{GW}(k)$ only when $Y$ is $\mathbb{A}^{1}$-connected. Therefore, one might expect $\mathbb{A}^{1}$-connectedness to be a necessary condition for symmetrisability. However, when $k$ is a number field, there exist smooth projective curves of genus $\geq 1$ that are not $\mathbb{A}^{1}$-connected. It is therefore not true that over a number field $\mathrm{K}_{0}(\mathrm{Sym}_{k})$ is given by the Grothendieck group of (geometrically)-$\mathbb{A}^{1}$-connected varieties.